# QUB-Cirdan at “Discharge Me!”: Zero shot discharge letter generation by open-source LLM

### v1

@@ Abstract
Rui Guo1,2 Greg Farnan2 Niall McLaughlin1 Barry Devereux1 1 Queen’s University Belfast 2 Cirdan Correspondence: rui.guo@cirdan.com
@@ 1 Introduction
A discharge letter, or a discharge summary, is a critical document summarizing a patient’s hospital visit from admission to discharge, serving as a bridge between hospital care and follow-up with outpatient providers. Among its several sections, the “Brief Hospital Course” outlines the patient’s treatment and progress during the hospital stay, typically using clinical jargon that is best understood by healthcare professionals. Conversely, the “Discharge Instructions” are designed to guide patients and their caregivers once they leave the hospital, using layman’s language to clearly explain follow-up care, medication regimens, and lifestyle recommendations.
Large Language Models (LLMs) offer a promising solution for automating medical documentation due to their ability to understand and generate human-like text (Singhal et al., 2023a; Zhang et al., 2023). Unlike traditional extractive summarization El-Kassas et al. (2021), which predominantly involves concatenating snippets from existing texts, LLMs can enhance summarization by integrating both extractive and abstractive techniques. This has been applied to progress note summarization (Gao et al., 2022; Liu et al., 2023), which is similar to this Codabench challenge. With both proprietary LLMs such as ChatGPT (OpenAI, 2024) and open-source LLMs such as Llama3 (AI@Meta, 2024), the potential for creating accessible medical summaries is significant.
@@
Crafting specialized templates for the “Brief Hospital Course” and “Discharge Instructions” sections, with carefully designed prompts to ensure that the generated text is medically reliable and stylistically consistent with the training dataset.
@@ 2 Related Work
The application of foundation models, pre-trained on billions of tokens from diverse data sources, is increasingly prevalent in healthcare (He et al., 2024). These models are pivotal in various domains, such as diagnosis generation (Gao et al., 2023b) and medical image analysis (Zhang et al., 2024). Within clinical text processing, large language models (LLMs) are employed for tasks including summarization (Van Veen et al., 2023; Gao et al., 2023a) and answering medical questions (Singhal et al., 2023b). Specifically, the “Discharge me!” challenge involves condensing extensive medical records into succinct discharge letters while retaining all critical information, making LLMs ideally suited for this task.
@@
The process of pre-training and fine-tuning LLMs typically requires GPUs with significant memory capacities (often exceeding 16GB). Fine-tuning can take several days, even when using Parameter-Efficient Fine Tuning (PEFT) methods like LoRA (Hu et al., 2021). However, if provided with the appropriate context and instructions, modern LLMs can exhibit strong performance without additional fine-tuning. For instance, Almanac (Zakka et al., 2024) enhances its output by retrieving clinical question-related knowledge from curated sources, a technique known as Retrieval-augmented Generation (RAG) (Gao et al., 2023c). Additionally, Medagents (Tang et al., 2023) demonstrates that a zero-shot method, which deconstructs the question into distinct steps and assigns specific prompts and roles to the LLM for each stage, can achieve competitive results compared to more traditional few-shot approaches.
@@ 3 Methods
While GPT-4/3.5 models generally outperform open-source models such as Llama2 in understanding EHR data (Liu et al., 2024), the rules of this challenge discourage the use of proprietary model APIs (e.g., OpenAI’s GPT-4). Consequently, we resorted to the state-of-the-art (SOTA) open-source model, Llama3 (AI@Meta, 2024). Our approach leverages the full text from the “text” field in the provided discharge.csv file, alongside aggregated fields from other MIMIC-IV tables, including patient information, diagnoses, and transfer history. We meticulously curated a template for each target section and designed prompts to guide the LLM in generating the required sections. In addition to our final approach, we documented several other zero-shot methods for target section generation and various approaches to predict the target section’s word count, although these were not adopted in our final solution.
@@ 3.1 Dataset Exploration
The dataset for this challenge is derived from MIMIC-IV’s submodules, MIMIC-IV-Note (Johnson et al., 2023c) and MIMIC-IV-ED (Johnson et al., 2023a). All patients included have visited the Emergency Department (ED), and the final target sections, “Brief Hospital Course” and “Discharge Instructions”, are extracted from their discharge letters. Since patients can be admitted to the hospital after their initial ED visit, we also explored other tables from the MIMIC-IV hosp and ICU modules (Johnson et al., 2023b) to provide a comprehensive view of the patient’s hospital stay beyond the ED information.
Due to limited context length, we could not simply pass all available information into the LLM. Therefore, we ranked all sections of the discharge letter to select a subset of the information. We segmented the discharge letter’s “text” column from discharge.csv using regex and a template of keywords for different sections, as shown in the Section column of Table 1. Besides the information from the “text” column, we also aggregated “Patient Admissions” information, including gender, race, age (calculated), “Diagnoses” (throughout the patient stay), and “Transfer Summary” from other MIMIC-IV tables. We then calculated the average ranking of the metric score for each section relative to the target sections, using the provided evaluation metrics (Xu et al., 2024). Each section was compared to the target sections, “Brief Hospital Course” (BHC) and “Discharge Instructions” (DI), with higher-ranking sections being more related to the target sections. Table 1 shows that “History of Present Illness” is most related to the BHC section, followed by imaging results, physical exams, past medical history, and diagnoses. BHC is most related to DI, followed by sections related to BHC.
@@ 3.3 Target Section Structure Template and Prompt Creation
The template includes several optional sections that are not included in all the examples. The template will be fed to the prompt below as the “structure” variable. The prompt for BHC is:
@@ 4 Results
Baseline with Random Shuffling: We shuffled the “hadm_id” column, a unique identifier for each patient’s discharge letter, assigning a random target section to each “hadm_id”.
@@
Baseline with RAG Retrieval: We used the retrieved target sections directly.
@@
Table 2 presents the evaluation metrics from the Codabench platform (Xu et al., 2024), including BLEU, ROUGE-1, ROUGE-2, ROUGE-L, BERT, METEOR, Align, and MEDCON. The random shuffle yielded the lowest scores across all metrics, indicating poor performance. Using the retrieved target section directly resulted in the highest BLEU score. The fixed word count approach achieved higher Align and MEDCON scores than the retrieved target section but had lower scores for other metrics. Our proposed method, which combines the retrieved word count and structured template, achieved the highest scores across all metrics except BLEU. The lower BLEU score for the proposed method is due to BLEU’s heavy penalty for deviations from exact wording. In contrast, the higher ROUGE scores indicate our method effectively captures the essential content, even with varied wording. We also measured the generation time for each section. The average time to generate one BHC was 16.67 seconds, and one DI was 16 seconds.
@@ 6 Conclusion
In this paper, we present a resource-friendly approach to automating the generation of the “Brief Hospital Course” and “Discharge Instructions” sections in discharge letters using the Llama3 8B quantized model. Our zero-shot template-based method and Retrieval-Augmented Generation produce high-quality, contextually appropriate summaries. However, we observe a lower BLEU score due to the different wording between the method’s result and the target sections. Ensuring the reliability and accuracy of generated content remains a significant challenge. Future work will focus on enhancing model reasoning capabilities, improving dynamic template extraction, and integrating robust validation mechanisms to verify medical accuracy.
@@ 7 Ethical Statement
All the data used in the experiments are downloaded from the PhysioNet after completing the required CITI training and credentialing process. Beyond the general potential ethical considerations of using LLMs to automatically process and generate clinical text (including issues of bias, fairness, transparency and accountability), there are no specific ethical issues raised by the particular methodologies or data presented in this research.

### v2

@@ Abstract
Rui Guo1,2 ††thanks: rui.guo@cirdan.com Greg Farnan2 ††thanks: greg.farnan@cirdan.com Niall McLaughlin1 ††thanks: n.mclaughlin@qub.ac.uk Barry Devereux1 ††thanks: b.devereux@qub.ac.uk 1 Queen’s University Belfast 2 Cirdan
@@ 1 Introduction
A discharge letter, or a discharge summary, is a critical document summarizing a patient’s hospital visit from admission to discharge, serving as a bridge between hospital care and follow-up with outpatient providers. Among its several sections, the “Brief Hospital Course” outlines the patient’s treatment and progress during the hospital stay, typically using clinical jargon best understood by healthcare professionals. Conversely, the “Discharge Instructions” are designed to guide patients and their caregivers once they leave the hospital, using layman’s language to clearly explain follow-up care, medication regimens, and lifestyle recommendations.
Large Language Models (LLMs) offer a promising solution for automating medical documentation due to their ability to understand and generate human-like text (Singhal et al., 2023a; Zhang et al., 2023). Unlike traditional extractive summarization El-Kassas et al. (2021), which predominantly involves concatenating snippets from existing texts, LLMs can enhance summarization by integrating both extractive and abstractive techniques. This has been applied to progress note summarization (Gao et al., 2022; Liu et al., 2023), similar to this Codabench challenge. With both proprietary LLMs such as ChatGPT (OpenAI, 2024) and open-source LLMs such as Llama3 (AI@Meta, 2024), the potential for creating accessible medical summaries is significant.
@@
Crafting specialized templates for the “Brief Hospital Course” and “Discharge Instructions” sections, with carefully designed prompts to ensure the generated text is medically reliable and stylistically consistent with the training dataset.
@@ 2 Related Work
The application of foundation models, pre-trained on billions of tokens from diverse data sources, is increasingly prevalent in healthcare (He et al., 2024). These models are pivotal in various domains, such as diagnosis generation (Gao et al., 2023b) and medical image analysis (Zhang et al., 2024). Within clinical text processing, large language models (LLMs) are employed for tasks including summarization (Van Veen et al., 2023; Gao et al., 2023a) and answering medical questions (Singhal et al., 2023b). Specifically, the “Discharge me!” challenge involves condensing extensive medical records into succinct discharge letters while retaining all critical information, making LLMs suited for this task.
@@
Pre-training and fine-tuning LLMs require GPUs with significant memory capacities (often exceeding 16GB). Fine-tuning can take several days, even using Parameter-Efficient Fine Tuning (PEFT) methods like LoRA (Hu et al., 2021). However, modern LLMs can exhibit strong performance without additional fine-tuning if provided with the appropriate context and instructions. For instance, Almanac (Zakka et al., 2024) enhances its output by retrieving clinical question-related knowledge from curated sources, a technique known as Retrieval-augmented Generation (RAG) (Gao et al., 2023c). Additionally, Medagents (Tang et al., 2023) demonstrates that a zero-shot method, which deconstructs the question into distinct steps and assigns specific prompts and roles to the LLM for each stage, can achieve competitive results compared to more traditional few-shot approaches.
@@ 3 Methods
While GPT-4/3.5 models generally outperform open-source models such as Llama2 in understanding EHR data (Liu et al., 2024), the rules of this challenge discourage the use of proprietary model APIs (e.g., OpenAI’s GPT-4). Consequently, we resorted to the state-of-the-art (SOTA) open-source model, Llama3 (AI@Meta, 2024). Our approach leverages the full text from the “text” field in the provided discharge.csv file, alongside aggregated fields from other MIMIC-IV tables, including patient information, diagnoses, and transfer history. We meticulously curated a template for each target section and designed prompts to guide the LLM in generating the required sections. In addition to our final approach, we documented several other zero-shot methods for target section generation and various approaches to predict the target section’s word count. However, these were not adopted in our final solution.
@@ 3.1 Dataset Exploration
The dataset for this challenge is derived from MIMIC-IV’s submodules, MIMIC-IV-Note (Johnson et al., 2023c) and MIMIC-IV-ED (Johnson et al., 2023a). All patients have visited the Emergency Department (ED), and the final target sections, “Brief Hospital Course” and “Discharge Instructions”, are extracted from their discharge letters. Since patients can be admitted to the hospital after their initial ED visit, we also explored other tables from the MIMIC-IV hosp and ICU modules (Johnson et al., 2023b) to provide a comprehensive view of the patient’s hospital stay beyond the ED information.
Due to limited context length, we could not simply pass all available information into the LLM. Therefore, we ranked all sections of the discharge letter to select a subset of the information. We segmented the discharge letter’s “text” column from discharge.csv using regex and a template of keywords for different sections, as shown in the Section column of Table 1. Besides the information from the “text” column, we aggregated “Patient Admissions” information, including gender, race, age (calculated), “Diagnoses” (throughout the patient stay), and “Transfer Summary” from other MIMIC-IV tables. Since we compiled the patient’s diagnoses and transfer summary for the entire hospital stay using other MIMIC-IV tables rather than just the Emergency Department (ED) stay, we did not use the tables in the ED module, such as triage, edstays, and diagnosis, as they only cover part of the patient’s stay. The content of “radiology” will be set to the content of the section “Imaging” if the “Imaging” section is empty in the discharge letter. We then calculated the average ranking of the metric score for each section relative to the target sections, using the provided evaluation metrics, including BLEU-4 (Papineni et al., 2002), ROUGE-1/2/L (Lin, 2004), BERTScore (Zhang et al., 2019), Meteor (Banerjee and Lavie, 2005), AlignScore (Zha et al., 2023), and MEDCON (Yim et al., 2023). Each section was compared to the target sections, “Brief Hospital Course” (BHC) and “Discharge Instructions” (DI), with higher-ranking sections being more related to the target sections. Table 1 shows that “History of Present Illness” is most related to the BHC section, followed by imaging results, physical exams, past medical history, and diagnoses. BHC is most related to DI, followed by sections related to BHC.
@@ 3.3 Target Section Structure Template and Prompt Creation
The template includes several optional sections not included in all the examples. The template will be fed to the prompt below as the “structure” variable. The prompt for BHC is:
@@ 4 Results
Baseline with Random Shuffling: We shuffled the “hadm_id” column, a unique identifier for each patient’s discharge letter, assigning a random target section to each “hadm_id”. This random selection comes from the same distribution as the training data but without the actual content of the input text.
@@
Baseline with RAG Retrieval: We used the retrieved target sections directly. This result can be similar to the target, but the details can differ from the real input.
@@
Table 2 presents the evaluation metrics from the Codabench platform (Xu et al., 2024), including BLEU-4 (Papineni et al., 2002), ROUGE-1/2/L (Lin, 2004), BERTScore (Zhang et al., 2019), Meteor (Banerjee and Lavie, 2005), AlignScore (Zha et al., 2023), and MEDCON (Yim et al., 2023). The random shuffle yielded the lowest scores across all metrics, indicating poor performance. Using the retrieved target section directly resulted in the highest BLEU score. The fixed word count approach achieved higher Align and MEDCON scores than the retrieved target section but had lower scores for other metrics. Our proposed method, which combines the retrieved word count and structured template, achieved the highest scores across all metrics except BLEU. The lower BLEU score for the proposed method is due to BLEU’s heavy penalty for deviations from exact wording. In contrast, the higher ROUGE scores indicate our method effectively captures the essential content, even with varied wording. We also measured the generation time for each section. The average time to generate one BHC was 16.67 seconds, and one DI was 16 seconds.
@@ 6 Conclusion
In this paper, we present a resource-friendly approach to automating the generation of the “Brief Hospital Course” and “Discharge Instructions” sections in discharge letters using the Llama3 8B quantized model. Our zero-shot template-based method and Retrieval-Augmented Generation produce high-quality, contextually appropriate summaries. However, we observe a lower BLEU score due to the different wording between the method’s result and the target sections. Ensuring the reliability and accuracy of generated content remains a significant challenge. Future work will focus on enhancing model reasoning capabilities, improving dynamic template extraction, and integrating robust validation mechanisms to verify medical accuracy. The code for this work is shared on https://github.com/ruiguo-bio/discharge_me, covering aggregating additional tables, segmentation of the discharge letters, RAG for the two target sections, and the random forest classifier for the target section words prediction.
7 Limitations and Future Work
1.
We would like to perform a more thorough evaluation to ensure that the model’s generated content is clinically relevant and does not include false or harmful information. This evaluation could be extended to understanding the strengths and weaknesses of language models for the challenge task.
2.
We create a template by sampling target sections with word counts close to the median. However, the length and structure of real target sections can vary significantly from our template. Our approach could be improved by predicting the target word count more precisely or by sampling different templates depending on the word count.
3.
We would like to test a wider range of language models and thoroughly compare different methods of providing relevant context to the language model, including different methods of Retrieval-Augmented Generation (RAG) and prompt engineering.
@@ 8 Ethical Statement
All the data used in the experiments are downloaded from the PhysioNet after completing the required CITI training and credentialing process. Beyond the general potential ethical considerations of using LLMs to automatically process and generate clinical text (including bias, fairness, transparency and accountability), there are no specific ethical issues raised by the particular methodologies or data presented in this research.